When two births are randomly​ selected, the sample space for genders is​ bb, bg,​ gb, and gg. Assume that those four outcomes are equally likely. Construct a table that describes the sampling distribution of the sample proportion of girls from two births. Does the mean of the sample proportions equal the proportion of girls in two​ births? Does the result suggest that a sample proportion is an unbiased estimator of a population​ proportion? For the entire​ population, assume the probability of having a boy is 1/2
​, the probability of having a girl is 1/2
​, and this is not affected by how many boys or girls have previously been born.
I have no clue how or what to do here.

Answer by Boreal(15235)

BB proportion 1/4
BG proportion 1/4
GB proportion 1.4
GG proportion 1/4
E(X) where x is number of girls expected from 2 births. It is the sum of all the random variables x*p(x)
=0*1/4=1/(1/4)+1(1/4)+2 (1/4)=1
So the mean number of girls in 2 births is 1.
This suggests that the sample proportion is an unbiased estimate of the population proportion, because the expected value is the same as the population proportion (1/2) and 1 girl in 2 pregnancies.